biggestbeefsteak:

im confused on how to answer sin (3x) = 1/2 , i know how to answer it but how am i supposed to know when to stop adding the period to the two solutions if i want to find all the solutions from 0 to 2pi

mintyy:

To know when to stop adding the period, you continue adding the period to your base solutions until the resulting x value is greater than or equal to \(2\pi \). The key is to find all possible solutions for the inner angle first, within an expanded range, and then solve for \(x\) and check if the result falls within the required \([0,2\pi )\) interval. 

biggestbeefsteak:

but how do you check that im confused because i got 29pi over 18 but how do u check that its smaller or bigger

Luigi0210:

You could use a calculator to check the exact values. For example: \(\large 2 pi \approx 6.2832 \) and \(\large \frac{29 }{18}\pi \approx 5.061 \) Which is within range of 2\(\pi \)

jhonyy9:

- so like a first step maybe good than you know that sin 30° = 1/2

Calculusexpert:

Product Function: sin(x) · cos(x) Derivative: cos(x) 2 - sin(x) 2 Calculate your derivatives at: https://www.derivativecalculus.com/